OpenStudy (anonymous):

if 1000 is decreased by 85% what is the new amount?

OpenStudy (tkhunny):

What is 85% of 1000?

OpenStudy (anonymous):

You would find 85% of 1000 (1000 x 0.85) and subtract that amount from 1000.

geerky42 (geerky42):

I'm sure it's just 1000 x 0.85. You don't have to subtract that from 1000 @Broadwaybaby

OpenStudy (anonymous):

oh yeah! I read the question wrong. @geerky42

OpenStudy (tkhunny):

Decreased TO 85% 1000 * 0.85 = 850 Decreased BY 85% 1000 - 1000 * 0.85 = 1000(1 - 0.85) = 1000 * 0.15 = 150